OpenStudy (anonymous):

alright here's a toughie.......Some students make necklaces and bracelets in their spare time and sell all that they make. They can make at most 40 bracelets and at most 80 necklaces each week. Every week, they have an available 10,000 grams of metal. It takes 50 grams of metal to make a necklace and 200 grams of metal to make a bracelet. The profit on each necklace is $3.25 and the profit on each bracelet is $3.50. How many of each should they make? how much profit can they make?

OpenStudy (anonymous):

1.02

OpenStudy (anonymous):

and that means ?

OpenStudy (anonymous):

That's how much profit.

OpenStudy (anonymous):

Most likely...I didn't solve it but you really can't make 1.02 of a bracelet.

OpenStudy (anonymous):

for each bracelet

OpenStudy (anonymous):

yeaa i'm confused..

OpenStudy (anonymous):

is this one of these annoying "linear programming" problems?

OpenStudy (anonymous):

maximize \(3.25x+3.50y\) subject to the condition that \(20x+200y\leq 10,000\)

OpenStudy (anonymous):

but why is it 20x?

OpenStudy (anonymous):

and also \(x\leq 40,y\leq 80\) like that i think

OpenStudy (anonymous):

because i made a typo, it is \(50x+200y\leq 10,000\)

OpenStudy (anonymous):

lol okay well i got all those equations except the 3.25x+3.50y. what does that equation equal?

OpenStudy (anonymous):

the profit if you make \(x\) bracelets and \(y\) necklaces

OpenStudy (anonymous):

that is what you are trying to maximize i think you check the corners if i am not mistaken

OpenStudy (anonymous):

oh i see! thank you!

OpenStudy (anonymous):

yw hope you know what to do from here because i forget well sort of. i know you need to check the corners, where the lines intersect and see which one gives you the biggest output for \(3.25x+3.50y\)

OpenStudy (anonymous):

oh yes. i know how to do it, i was just making sure before i finished